Define f(x) = sinx on [0, 2pi]. Find two increasing functions h and g for which f = h — g on 

[0, 2pi]. 


1
Expert's answer
2021-02-24T07:41:04-0500

We will search the functions "h,g:[0;2\pi]\to\mathbb{R}" which are "\mathcal{C}^1". We have "f'=(h-g)'=h'-g'". The functions "h,g" are strictly increasing if "h', g'>0" on "[0;2\pi]". Thus we need to find two positive continuous functions "h',g'" such that "h'-g'=\cos x". For "h'" we need to take something that is always strictly bigger than "\cos x", for example, "h'=2=const". In this case "g'=2-\cos x". Both are always positive. Now by finding their primitives "h=2x+C, g=2x-\sin(x)+C", we see that the difference of such two functions is clearly "\sin(x)" and they both are strictly increasing, as their derivatives are positive.